OpenStudy (anonymous):

substitution method 4x + y = 16 7x +3y = 9 @Kainui

OpenStudy (anonymous):

so far I have these steps completed... 4x + y =16 4x + y = 16 + 4x 7x 3(16 + 4) = 9 7x + 48 + -12) = 9

OpenStudy (anonymous):

@dan815 @TheSmartOne

OpenStudy (aaronandyson):

umm?

OpenStudy (aaronandyson):

4x + y = 16 7x +3y = 9 do this : 4x + y = 16 therefore, y = 16 - 4x now substitude

OpenStudy (anonymous):

I did that already...look at the response after my post

OpenStudy (aaronandyson):

that is 7x + 3(16 - 4x) = 9

OpenStudy (aaronandyson):

Followed so far?

OpenStudy (anonymous):

Do I take 48 from each side?

OpenStudy (aaronandyson):

wait

OpenStudy (aaronandyson):

i made a mistake

OpenStudy (anonymous):

I know after this step I have to do the distributive law...which means 7x + -12 = 5x...then I take 48 from each side which leaves me 31?

OpenStudy (anonymous):

The best way to do these, is to first look at the problem and isolate the variable that doesn't have a constant, right off the bat, we know that y= 16- 4x from the first equation. WE can use this, and sub in for y in the second equation: 7x + 3(16-4x) = 9 7x + 18 - 12x = 9 -5x + 18 = 9 -5x = -9 x = 9/5 Now sub x in the first equation to solve for y. 4(9/5) + y = 16 (36/ 5) + y = 16 y = 16 - 36 / 5 y = 8.8

OpenStudy (anonymous):

Let me know if you need a further explanation.

OpenStudy (anonymous):

its 48 not 18

OpenStudy (anonymous):

Woah, I made a huge mistake. 3*16 is not 18.. Here's a redo. 7x + 3(16-4x) = 9 7x + 48 - 12x = 9 -5x = -39 x = 39/5

OpenStudy (anonymous):

Sub in the second, we get 4(39/5) + y = 16 y = -15.2

OpenStudy (love_ranaa):

x= 7.8 y= -15.2 :)